WHAT IS THE hybridization of SO4(2-)

Grade:12th Pass

4 Answers

bhaumik shah
12 years ago

a simple metod valence electron in s are 6 and if nigbouring atoms are divalent or trivalent then dont count them and add the charge

6+2/2=4 so hybridisation is sp3 as there are 4 sp3 hybridised orbitals

Sanjeev Malik IIT BHU
12 years ago

S has 6 valence electrons . with divalent O it has S=O type bonding thus taking 2 elec. for each this type bonding.

2- charge on  S act as lone pair i.e 2 elec for S again used for 4th O. thus it has 4 bondings total giving sp3 hybridisation.

approve by clicking yes below

BABLU KUMAR
12 years ago

sp3 because it has 4 sigma bonds and no lone pair.so the total no of sigma bond + LP IS 4 .SO IT has 4 sp3 orbitals.

Trick:-> if sigma bond + lp is 2 then sp.

if sigma bond + lp is 3 then sp 2.

if sigma bond + lp is 4 then sp 3. and so on.

thanks pls approve it...............

Bhawna Singh
2 years ago
Formula to find hybrid orbitals = 1/2 ( V + M - C + A ) , where V = valence electron in central atom , M = Monovalent atoms in surroundings , C = Cationic Charge , A = Anionic Charge ..... Formula to find lone pair = No. Of hybrid orbitals - surrounding atoms ....... So here we have to find the Hybridisation of So42- as we only count no. Of monovalent atoms to find hybrid orbitals so oxygen would not be considered as it is not monovalent  now configuration of S is 2, 8, 6 we would be taking 6 as valence electron  ....  Hybrid orbitals = 1/2 ( 6 - (-2)) = 4 that means it's hybridization would be sp3... Now lone pair = 4 - no.of oxygen atom = 4 - 4 = 0 so 0 lone pair .... that shows that it's geometry would be tetrahedral...... That was my answer thank you so much I hope you will understand it now